the measure of an angle is 5 less than 4 times the measure of its supplement.
find the measures of both angles?

Answer by HyperBrain(694)

Let x=angle
Then 180-x=supplement of the angle
---
We have the equation:
x=4(180-x)-5
x=720-4x-5
x=715-4x
5x=715
x=143 degrees